# Kernel Density Steering: Inference-Time Scaling via Mode Seeking for Image Restoration

### v1

@@ A.1.4 Integration with Standard Diffusion Samplers
4:    $\mathbf{E}_{\theta,t}\leftarrow\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,\varnothing)+w(\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,{\bm{c}})-\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,\varnothing))$triangleright Predict ensemble noise with CFG
5:    $\hat{{\bm{Z}}}_{0|t}^{\text{pred}}\leftarrow({\bm{Z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{E}_{\theta,t})/\sqrt{\bar{\alpha}_{t}}$triangleright Predict $\mathbf{x}_{0}$ ensemble
6:    $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$triangleright Same as Step 5-15 in Algorithm 2
7:    $\mathbf{E}^{\prime}_{t}\leftarrow({\bm{Z}}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}})/\sqrt{1-\bar{\alpha}_{t}}$triangleright Update direction based on KDS-guided $\mathbf{x}_{0}$
8:    ${\bm{Z}}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}+\sqrt{1-\bar{\alpha}_{t-1}}\mathbf{E}^{\prime}_{t}$triangleright DDIM step with KDS-guided $\mathbf{x}_{0}$
@@
10:return $\hat{{\bm{Z}}}_{0}^{\text{KDS}}$triangleright Return KDS-guided result
@@
2:${\bm{Z}}_{T}\sim\mathcal{N}(0,\mathbf{I})$ (ensemble of $N$ samples) $\triangleright$ Init noise ensemble
3:$\tilde{\bm{Z}}_{t_{0}}\leftarrow{\bm{Z}}_{T}$.
@@
2:${\bm{Z}}_{T}\sim\mathcal{N}(0,\mathbf{I})$ (ensemble of $N$ samples) $\triangleright$ Init noise ensemble
3:$\tilde{\bm{Z}}_{t_{0}}\leftarrow{\bm{Z}}_{T}$.
@@
8:     $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$triangleright Same as Step 5-15 in Algorithm 2
@@
11:     $\hat{{\bm{U}}}_{0|s}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{U}}}_{0|s}^{\text{pred}})$triangleright Same as Step 5-15 in Algorithm 2
@@ A.1.6 Additional comparison with Best-of-N approaches:
While inference-time scaling allows for generating multiple candidate solutions, particularly useful for low-quality inputs, selecting the optimal one from an N-particle ensemble remains a challenge. In our main paper, we didn’t cover the full scope of this experiment. Here, we expand on that by including more metrics. As Table 8 now demonstrates, using non-reference metrics like LIQE [64] or ClipiQA to pick the best particle from an N-selection (i.e., ”best LIQE best of N” or ”best CLIPIQA best of N”) results in significantly lower performance compared to our Kernel Density Steering (KDS) method. This shows that, despite comparable computational costs, traditional post-sampling selection methods don’t achieve the same performance level as KDS. As shown in Figure 8, KDS method acheve the most stable performance compared with both BoN baselines, which suffers from the artifacts which confused the non-reference metrics.
@@
A.2 Experiment details: Image Inpainting
Datasets We generated our inpainting test set by center corpped $30\%$ square region of each image. We generate used first 1,000 images from ImageNet testset.
Hyperparameters: Similar to real-world SR settings, we fixed the patch size to 1, bandwidth $h$ to 0.3, steering strength $\delta_{t}$ same as introduced in (Eq: 10).
A.3 Additional Visual Results

### v2

@@ A.1.4 Integration with Standard Diffusion Samplers
4:  $\mathbf{E}_{\theta,t}\leftarrow\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,\varnothing)+w(\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,{\bm{c}})-\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,\varnothing))$ $\triangleright$ Predict ensemble noise with CFG
5:  $\hat{{\bm{Z}}}_{0|t}^{\text{pred}}\leftarrow({\bm{Z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{E}_{\theta,t})/\sqrt{\bar{\alpha}_{t}}$ $\triangleright$ Predict $\mathbf{x}_{0}$ ensemble
6:  $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$ $\triangleright$ Same as Step 5-15 in Algorithm 2
7:  $\mathbf{E}^{\prime}_{t}\leftarrow({\bm{Z}}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}})/\sqrt{1-\bar{\alpha}_{t}}$ $\triangleright$ Update direction based on KDS-guided $\mathbf{x}_{0}$
8:  ${\bm{Z}}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}+\sqrt{1-\bar{\alpha}_{t-1}}\mathbf{E}^{\prime}_{t}$ $\triangleright$ DDIM step with KDS-guided $\mathbf{x}_{0}$
@@
10:return $\hat{{\bm{Z}}}_{0}^{\text{KDS}}$ $\triangleright$ Return KDS-guided result
@@
[…]
@@
[…]
@@
8:  $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$ $\triangleright$ Same as Step 5-15 in Algorithm 2
@@
11:  $\hat{{\bm{U}}}_{0|s}^{\text{KDS}}\leftarrow\textbf{Patch-wise KDS}(\hat{{\bm{U}}}_{0|s}^{\text{pred}})$ $\triangleright$ Same as Step 5-15 in Algorithm 2
@@ A.1.6 Additional comparison with Best-of-N approaches:
While inference-time scaling allows for generating multiple candidate solutions, particularly useful for low-quality inputs, selecting the optimal one from an N-particle ensemble remains a challenge. In our main paper, we didn’t cover the full scope of this experiment. Here, we expand on that by including more metrics. As Table 9 now demonstrates, using non-reference metrics like LIQE [65] or ClipiQA to pick the best particle from an N-selection (i.e., ”best LIQE best of N” or ”best CLIPIQA best of N”) results in significantly lower performance compared to our Kernel Density Steering (KDS) method. This shows that, despite comparable computational costs, traditional post-sampling selection methods don’t achieve the same performance level as KDS. As shown in Figure 8, KDS method achieves the most stable performance compared with both BoN baselines, which suffers from the artifacts which confused the non-reference metrics.
@@
Method PSNR SSIM LPIPS ($\downarrow$) DISTS ($\downarrow$) LIQE CLIPIQA Time(s) Memory
DDIM 23.21 0.610 0.370 0.250 4.046 0.689 7.9s 8.0Gb
+ BoN (LIQE) 23.72 0.622 0.361 0.246 4.351 0.741 27.3s 10.4Gb
+ BoN (CLIPIQA) 23.01 0.592 0.382 0.247 4.187 0.774 28.9s 10.4Gb
+ KDS 24.39 0.669 0.339 0.245 3.819 0.692 28.4s 10.4Gb
A.1.7 KDS with post-hoc Best-of-N Selection:
We also investigate replacing the standard KDS final particle selection (Equation 10) with a post-hoc BoN selection. This involves first guiding the entire ensemble using KDS and then applying a non-reference metric (CLIPIQA) to select the best particle from the guided ensemble.
As shown in Table 10, this combination further validates our approach. KDS first enhances the overall quality of the ensemble, making the subsequent BoN selection more reliable. Without KDS, BoN selects from a noisy set of candidates and favors samples with artifacts that score well. With KDS, the candidates are already steered toward a high-fidelity consensus, providing a cleaner, more faithful set for the BoN selector. This results in a better perceptual–distortion trade-off than using BoN alone.
In summary, KDS provides a principled, task-aligned alternative to BoN by enforcing ensemble consistency. It improves performance across the board and can even strengthen BoN when the two are combined.
Table 10: Comparison of BoN and BoN + KDS Methods
@@ A.1.7 KDS with post-hoc Best-of-N Selection:
+ KDS+ BoN (CLIPIQA) 23.69 0.636 0.359 0.243 4.161 0.770
A.2 Additional Ablation Studies
This section provides the full empirical results for the ablation studies and comparisons mentioned in the main paper, including the full hyperparameter sweep for mode-seeking baselines and detailed ablations on KDS’s components.
A.2.1 Mode-Seeking Method Parameter Sweep
Table 11 provides the full parameter sweep for HDS [66] and DGS [67] on the RealSR dataset, supporting the analysis in Section 5.
Table 11: Full parameter sweep for HDS [66] and DGS [67] on RealSR.
Method PSNR SSIM LPIPS ($\downarrow$) DISTS ($\downarrow$) CLIPIQA
DiffBIR (Baseline) 23.21 0.610 0.370 0.250 0.689
+ HDS (t=20) 23.95 0.614 0.382 0.254 0.651
+ HDS (t=50) 24.33 0.633 0.385 0.266 0.650
+ HDS (t=100) 24.56 0.646 0.386 0.274 0.647
+ HDS (t=200) 25.00 0.665 0.378 0.288 0.567
+ HDS (t=500) 25.79 0.679 0.377 0.309 0.295
+ DGS (q=0.1) 22.57 0.580 0.381 0.253 0.679
+ DGS (q=0.3) 22.95 0.598 0.375 0.252 0.674
+ DGS (q=0.5) 23.21 0.610 0.370 0.249 0.689
+ DGS (q=0.99) 24.05 0.651 0.355 0.246 0.687
+ DGS (q=0.999) 24.03 0.647 0.357 0.246 0.680
+ KDS (N=10) 24.39 0.669 0.339 0.244 0.692
A.2.2 Ablation on Guidance and Kernel Functions
Table 12 investigates the choice of kernel function and compares our mode-seeking approach to a simpler L2 guidance. The results show consistent performance between Gaussian and Laplacian kernels, demonstrating robustness. Both kernel-based methods outperform the L2 guidance baseline in perceptual metrics, validating our mode-seeking design.
Table 12: KDS with different guidance and kernels on RealSR dataset.
Method PSNR SSIM LPIPS ($\downarrow$) DISTS ($\downarrow$) CLIPIQA
DiffBIR 23.21 0.610 0.370 0.250 0.689
+ KDS (Gaussian) 24.39 0.669 0.339 0.244 0.692
+ KDS (Laplacian) 24.30 0.665 0.337 0.241 0.690
+ KDS (L2) 24.63 0.679 0.349 0.257 0.663
A.2.3 Ablation on Automatic Bandwidth Selection
Table 13 demonstrates KDS’s robustness to bandwidth selection. An automatic median heuristic strategy (auto bw) achieves performance nearly identical to the manually tuned bandwidth (bw=0.3), confirming KDS is not reliant on tedious tuning.
Table 13: Effect of automatic bandwidth selection for KDS.
Method PSNR SSIM LPIPS ($\downarrow$) DISTS ($\downarrow$) CLIPIQA
DiffBIR 23.21 0.610 0.370 0.250 0.689
+ KDS (bw=0.3) 24.39 0.669 0.339 0.244 0.692
+ KDS (auto bw) 24.24 0.661 0.338 0.237 0.691
A.2.4 Ablation on Patch Size and Particle Count
Table 14 explores the trade-off between patch size and particle count ($N$). As shown, increasing the number of particles (e.g., from $N=10$ to $N=40$) allows for the effective use of a larger patch size (from 1 to 4), which in turn enhances overall performance by better mitigating the curse of dimensionality.
Table 14: Ablation Study on patch size and particle count ($N$).
PSNR SSIM LPIPS $\downarrow$ FID $\downarrow$
DDIM 22.05 0.531 0.307 20.89
+ KDS (patch-size=1, $N=10$) 22.52 0.555 0.289 20.33
+ KDS (patch-size=4, $N=40$) 22.70 0.563 0.281 20.08
A.3 Experiment details: Image Inpainting
Datasets We generated our inpainting test set by center cropping $30\%$ square region of each image. We generate used first 1,000 images from ImageNet testset.
Hyperparameters: Similar to real-world SR settings, we fixed the patch size to 1, bandwidth $h$ to 0.3, steering strength $\delta_{t}$ same as introduced in (Eq: 11).
A.4 Additional Visual Results